OpenStudy (anonymous):

differentiate y = x^3(1+x)^3

OpenStudy (anonymous):

u = x^3 u' = 3x^2 v = (1+x)^3 v'=3(1+x)^2 y' = u'v + uv'

OpenStudy (anonymous):

use product rule

OpenStudy (anonymous):

so what would you get as an answer

OpenStudy (anonymous):

what answer did you get?

OpenStudy (anonymous):

combine all to y'

OpenStudy (anonymous):

i got 3x^2(1+x)^3 + 3x^3(1+x)^2

OpenStudy (anonymous):

thats true

OpenStudy (anonymous):

but the solutions at the end of the text book say that the answer is 3x^2(1+x)^2(1+2x)

OpenStudy (anonymous):

so ???

OpenStudy (anonymous):

You can do it the long way and evaluate \[(1+x)^3 = 1+3x+ 3x^2+ x^3 \] Then multiply it with \[x^3\] and simply differentiate

OpenStudy (anonymous):

oh ok thank you romero! =D

OpenStudy (anonymous):

I mean its tedious and there are other ways that are much easier but you can always do that. I recommend learning substitution. You will come upon problems where simply evaluating will not work.

OpenStudy (anonymous):

ohhhh like that, sorrry cant help much

OpenStudy (anonymous):

oh ok a BIG THANKS to all!